OpenStudy (anonymous):

Which expression is equivalent to 1 – cos^2 x? A. csc2 x B. sec2x C. sin2 x D. -sec2 x E. -sin2 x

OpenStudy (anonymous):

\[\sin^2x\]

OpenStudy (anonymous):

Simplify the expression: tan(-x)csc(x). Select one of the options below as your answer: A. cos x B. -cos x C. sin x D. sec x E. -sec x

OpenStudy (anonymous):

\[\frac{ -\sin x \times \csc x }{ \cos x }=\frac{ \sin x \times \frac{ 1 }{\sin x } }{ \cos x }=\frac{ -\sin x }{ \sin x \times \cos x }=\frac{ -1 }{ \cos x }=-\sec x\]